Dehner Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on direct labor-hours. The company based its predetermined overhead rate for the current year on the following data:





Total direct labor-hours 40,000
Total fixed manufacturing overhead cost $ 96,000
Variable manufacturing overhead per direct labor-hour $ 3.00

Recently, Job P951 was completed with the following characteristics:





Number of units in the job 20
Total direct labor-hours 100
Direct materials $ 755
Direct labor cost $ 4,000



The total job cost for Job P951 is closest to:
1
Expert's answer
2020-12-10T13:43:23-0500

Estimated total manufacturing overhead cost = $96,000 + ($3.00× 40,000) = $216,000.

Predetermined overhead rate = $216,000 ÷ 40,000 = $5.40.

Overhead applied to a particular job = $5.40×100 = $540.

Total cost of Job P951 = 755 + 4,000 + 540 = $ 5,295.